let log2=a log3 =b and log 5= c, what is log 16 log 0.00003 log 8/3 in terms of a, b , c? please help
Answer by nerdybill(7384)

let log2=a log3 =b and log 5= c, what is log 16 log 0.00003 log 8/3 in terms of a, b , c?
.
log(16)
= log(2^4)
= 4*log(2)
= 4a
.
log(0.00003)
= log(3/10^5)
= log(3) - log(10^5)
= log(3) - 5log(10)
= log(3) - 5log(2*5)
= log(3) - 5[log(2)+log(5)]
= b-5[a+c]
.
log 8/3
= log(8)-log(3)
= log(2^3)-log(3)
= 3log(2)-log(3)
= 3a-b